[figure: 9.7 cm]

SPIN DYNAMICS OF FERROMAGNETIC SPIN GLASS

P. RUSEK Institute of Physics, Wroclaw University of Technology, Wroclaw, PL

In the ferromagnetic spin glass (FSG) positions and directions of spins are randomly distributed, but the system has a finite spontaneous magnetization M whose value, at zero temperature, however, differs considerably from the nominal value. Recently the interest to FSG is rapidly growing because of similarity between FSG and ferromagneticlike phase in magnetic semiconductors- materials whose magnetic properties can be easily control-ed by the electronic parameters and vice versa. In FSG, in opposite to a conventional ferromagnet, the dynamically induced magnetic moment m essentially affects spin dynamics, because of the relatively small value of M. In a macroscopic description m can be treated as the spin of a vectorial field φ⃗ that describes rotation of spins with respect to a lattice. Since FSG is the low temperature phase the quantum effects are important in FSG. Starting with the phenomenological Lagrangian of the field φ⃗ in FSG we have derived the Hamiltonian of FSG and then canonically quantized the field φ⃗. It allows for the description of FSG spin dynamics at T ≠ 0, the inclusion of spin waves interaction, and correct inclusion of a dissipation in the system by taking into account the damping of spin waves. Using Green functions technique for T ≠ 0 the dynamical susceptibility of FSG and the cross section for inelastic neutron scattering has been calculated. Because of the quadratic dispersion law of spin waves in FSG the q vector and temperature dependence of those quantities are like the convectional ferromagnets.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : P. Rusek

Address for correspondence : Institute of Physics, Wroclaw University of Technology, Wroclaw, PL

Email address : rusekp@rainbow.if.pwr.wroc.pl